Solve the following equations and check your results 2y + 5/3 = 26/3 – y 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2y+5/3=26/3-y
⇒2y+y=26/3–5/3
⇒3y=(26–5)/3
⇒3y=21/3
⇒3y=7
⇒y=7/3
Putting the value of y in RHS and LHS we get,
⇒(2 × 7/3) + 5/3 = 26/3–7/3
⇒14/3 + 5/3 = 26/3–7/3
⇒(14+5)/3=(26-7)/3
⇒19/3 = 19/3
⇒LHS=RHS

Thanks